OpenStudy (anonymous):

3x+4=2x-5 Can someone help me please?

OpenStudy (anonymous):

keep x terms on one side and constants on the other side!

OpenStudy (anonymous):

Help with what, exactly?

OpenStudy (anonymous):

3x-2x=-5-4 gives out x=-9

OpenStudy (anonymous):

3x+2x=-4-5 5x=-9 5x/5=-9/5 x=-9/5?

OpenStudy (anonymous):

Ok, there is a problem, let's start at the beginning 3x+4=2x-5 Now subtract 2x from both sides.

OpenStudy (anonymous):

help

OpenStudy (anonymous):

See my post!

OpenStudy (anonymous):

1x+4=-5-4

OpenStudy (anonymous):

What's that?

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

or is it -6

OpenStudy (anonymous):

3x+4=2x-5 3x+4-4=2x-5-4 3x+2x=-6 5x=-6 5